# Large Language Models Might Not Care What You Are Saying: Prompt Format Beats Descriptions

### v1

@@ 2.7 Discussion
Above results show that LLMs benefit from our Ensemble prompts in most cases. However, the benefit comes from a proper format rather than the carefully designed meaningful descriptions (e.g., ”similar words” and ”similar syntax”). This demonstrates that LLMs might not necessarily care what users say in the descriptions but is more sensitive to the format of prompts. In other words, designing a proper prompt format would be more efficient than paying a lot of effort into looking for a perfect description.
In the following section, we apply the superior Ensemble format to more tasks to further verify its generalization.
@@ 3 Generalizing the New Ensemble Prompt Framework to More Tasks
Figure 5: Results on all nine datasets across three small-scale models. In the ”Date” subplot, the scores of Llama2 and Mistral under the Vanilla prompt are too low to be visible bars in the chart.
@@ 3.1 Experimental Setup
Datasets
We conduct a total of nine benchmarks, covering four types of tasks: commonsense QA, logical reasoning, arithmetic reasoning, and hallucination detection.
For commonsense QA, we adopt four datasets that either are widely used or focus on certain specific aspects of knowledge. The widely-used CSQA (Talmor et al. 2019) features commonsense questions about the world involving complex semantics that often require prior knowledge. StrategyQA (Geva et al. 2021) challenges models to infer implicit reasoning steps using a strategy to answer questions. Additionally, we choose two specialized evaluation sets from the BIG-bench (Srivastava et al. 2023): Date Understanding, which asks models to infer the date from a context, and Sports Understanding, which involves assessing the plausibility of sentences related to sports.
For logical reasoning task, we choose Logical Fallacy and Three Objects (a subset of Logical Deduction) from Big-bench (Srivastava et al. 2023). Logical Fallacy aims to test the model’s ability to identify whether there are fallacies in a given logical reasoning, and Three Objects requires the model to infer the order of a sequence of objects from a set of minimal conditions.
To explore the performance of ERR on math word problems, we adopt the following two datasets: GSM8K (Cobbe et al. 2021), which consists of high quality free-response grade school math problems, and AQuA (Ling et al. 2017), containing the algebraic word problems in the form of multiple-choice questions.
In addition, in order to explore whether ERR could alleviate LLMs’ hallucination, we choose Known Unknowns from Big-bench (Srivastava et al. 2023).
The details of splitting training set (example database) and test set are presented in supplementary materials.
Evaluation Metric
@@
Language Models
We evaluate three open source LLMs with 7B parameters, including Alpaca (Taori et al. 2023), Llama2 (Touvron et al. 2023b), and Mistral (Jiang et al. 2023). To assess the effect of ERR on more powerful models, we also evaluate GPT-3.5 (Ouyang et al. 2022). We use Llama-2-7b-chat-hf, Mistral-7B-Instruct-v0.2 and gpt-3.5-turbo-0125333https://openai.com/api/ for Llama2, Mistral and GPT-3.5 respectively.
Example Selection
@@
Prompts
We compare ERR with Vanilla across different datasets and LLMs. Given that these tasks usually involve reasoning, on which chain-of-thought (CoT) is commonly utilized (Wei et al. 2022), we experiment both without CoT (”w/o CoT”, which are identical to the original templates) and with CoT (”w/ CoT”). This allows us to examine both the orthogonality and compatibility with CoT of ERR , as well as assess its performance across various models and tasks. Specifically, we evaluate Vanilla (w/o CoT), Vanilla (w/ CoT), ERR (w/o CoT), and ERR (w/ CoT). Due to space constraints, examples of prompt templates discussed in this section are provided in the supplementary materials.
Figure 6: Results of all the four types of tasks on GPT-3.5.
@@ 3.2 Results of Small-scale Models
Results across all nine datasets and three small-scale models (Alpaca, Llama2 and Mistral) are illustrated in Figure 5.
The results demonstrate that ERR (w/ CoT), achieved by integrating CoT with our proposed prompt framework, either significantly outperforms or matches Vanilla (w/ CoT) in 25 out of 27 experiments (covering 9 datasets and 3 models). The exceptions are Alpaca on the Sports dataset and Mistral on the AQuA dataset, where ERR (w/ CoT) shows somewhat lower performance compared to Vanilla (w/ CoT). When CoT is not incorporated, ERR generally performs much better than or on par with Vanilla, except for the Logical Fallacy and Sports datasets with Llama2. In these cases, Llama2 frequently produces emojis or alternative questions under the ERR framework.
Surprisingly, ERR (w/o CoT) sometimes even surpasses Vanilla (w/ CoT), suggesting that the ERR framework alone can offer more improvements than CoT. This highlights the value of ERR and reaffirms that the format plays a crucial role in enhancing the model’s problem-solving capabilities.
In terms of models, the performance of ERR on Alpaca is far less impressive than on Llama2 and Mistral. This phenomenon also appears in GPT-3.5, which will be discussed in the next section. Another noteworthy observation is that while ERR outperforms Vanilla with Llama2 across most datasets, its performance on the Logical Fallacy and Sports datasets is notably poor. As mentioned earlier, Llama2 almost always responds with confused emojis for Logical Fallacy questions and outputs questions like ”plausible or implausible?” for Sports, leading to predominantly incorrect answers. Further investigation into these issues is left for future work.
@@
As shown in Figure 6, ERR performs similarly to Vanilla across every datasets using GPT-3.5. Although the ERR format does not bring significant improvement to these tasks with GPT-3.5 and Alpaca (as shown in Figure 5), the fact remains that the incorrect or meaningless example-level descriptions caused by random nouns do not have much negative impact on GPT-3.5, a sufficiently powerful model, or Alpaca, which has strong instruction-following capabilities. In some cases, it even slightly improves performance (e.g., ERR (w/ CoT) outperforms Vanilla (w/ CoT) on AQuA and Known Unknowns datasets). In other words, LLMs might not necessarily care what users actually say to describe the provided examples while they are more sensitive to the format of prompts, which is in line with our findings obtained from MT.
@@ 3.4 Discussion
Based on the experiments conducted on both small-scale and large-scale models, we can conclude that ERR is a simple yet practical and universal prompt framework. It can enhance problem-solving capabilities in small models and be applied to large models without the risk of performance degradation due to the meaningless noise within it. In other words, there might be less need to meticulously select examples or design detailed descriptions. Instead, you can uniformly and efficiently apply ERR to prompt engineering of various tasks with different models.
@@ 4 Related Work
Rubin, Herzig, and Berant (2022) suggest that LLMs’ ICL performance strongly depends on the selection of in-context examples. In consequence, many works have been trying to explore ways of selecting better in-context examples in recent years. Li et al. (2023) train a unified in-context example retriever across a wide range of tasks. Ye et al. (2023) select examples based on both relevance and diversity, with the help of determinantal point processes. Agrawal et al. (2023) ensure n-gram coverage to select better examples for MT. Kumar et al. (2023) train an in-context example scorer for MT based on several features. Tang, Wang, and Wu (2024) propose to combine both word-level and syntax-level coverage when selecting examples for MT.
@@
With the popularity of ICL, there have been numerous studies on analyzing the mechanism of ICL. One stream of these studies focuses on explaining the essence of ICL, relating ICL to gradient descent (Von Oswald et al. 2023), implicit Bayesian inference (Xie et al. 2022), induction heads completing token sequences based on similar context (Olsson et al. 2022), generation maintaining coherency (Sia and Duh 2023), creation of task vectors based on in-context examples (Hendel, Geva, and Globerson 2023), etc. The other stream focuses on the role of in-context examples, especially labels of these examples. Min et al. (2022) find that ground truth labels are not necessary and LLMs perform fairly well even with random labels, suggesting that LLMs mainly learn the label space, the input distribution and the overall task format from in-context examples. Wang et al. (2023) find that label words play the role of anchors that aggregating information of the whole examples and serve as a reference for LLMs’ final predictions. Wei et al. (2023) find that larger language models are able to override semantic priors and successfully learn from in-context examples with flipped labels or even semantically-unrelated labels.
@@ 5 Conclusion
In this work, we analyze the effect of descriptive instructions in prompts during ICL and propose an ensemble prompt framework describing the properties of in-context examples selected by different selection methods. Experimental results on MT indicate that while LLMs are sensitive to prompt formats, they might not necessarily care the actual meaning of the descriptions and the framework improves LLMs’ performance even with meaningless descriptions compared with the conventional prompt. We further apply the ensemble framework to four other NLP tasks and find that it achieves promising results, especially on small-scale models. These results suggest that rather than working hard on well-designed descriptions, making use of a proper prompt format would be more effective and efficient.

### v4

@@ 2.7 Discussion
Above results show that LLMs benefit from our Ensemble prompts in most cases. However, the benefit comes from a proper format rather than the meaningful descriptions (e.g., "similar words" and "similar syntax"). This demonstrates that LLMs might not care what users say in the descriptions but is more sensitive to the format of prompts. In other words, designing a proper prompt format would be more efficient than paying a lot of effort into looking for a perfect description.
In the next section, we apply Ensemble format to more tasks to further verify its generalizability.
@@ 3 Generalizing the New Ensemble Prompt Framework to More Tasks
Figure 6: Results on nine datasets across three small-scale models. In the "Date" subplot, the score of Mistral under the Vanilla prompt is too low to be a visible bar in the chart.
@@ 3.1 Experimental Setup
3.1.1 Datasets
We use a total of nine benchmarks, covering four task types: commonsense QA, logical reasoning, arithmetic reasoning, and hallucination detection.
For commonsense QA, we adopt four datasets. The widely-used CSQA Talmor et al. (2019) features commonsense questions about the world involving complex semantics requiring prior knowledge. StrategyQA Geva et al. (2021) challenges models to infer implicit reasoning steps using a strategy to answer questions. We also choose two specialized evaluation sets from BIG-bench Srivastava et al. (2023): Date Understanding, which asks models to infer the date from a context, and Sports Understanding, which involves assessing the plausibility of sentences related to sports.
For logical reasoning task, we choose Logical Fallacy and Three Objects (a subset of Logical Deduction) from Big-bench Srivastava et al. (2023). Logical Fallacy aims to test the model’s ability to identify whether there are fallacies in a given logical reasoning, and Three Objects requires the model to infer the order of a sequence of objects from a set of minimal conditions.
To explore the performance of ERR on math word problems, we adopt the following two datasets: GSM8K Cobbe et al. (2021), which consists of high quality free-response grade school math problems, and AQuA Ling et al. (2017), containing the algebraic word problems in the form of multiple-choice questions.
In addition, to explore whether ERR could alleviate LLMs’ hallucination, we choose Known Unknowns from Big-bench Srivastava et al. (2023).
Dataset Test Inputs
CSQA 1221
StrategyQA 1012
Date 365
Sports 996
Logical Fallacy 1012
Three Objects 296
Known Unknowns 42
GSM8K 1319
AQuA 254
Table 3: Number of test inputs for each dataset.
The number of test inputs for each dataset is listed in Table 3. Details of splitting training set (example database) and test set are in Appendix C.
3.1.2 Evaluation Metric
@@
3.1.3 Language Models
We evaluate three open source LLMs with around 7B parameters, including Alpaca Taori et al. (2023), Llama3 Grattafiori et al. (2024), and Mistral Jiang et al. (2023). To assess the effect of ERR on more powerful models, we also evaluate GPT-3.5 Ouyang et al. (2022). We use Llama-3.1-8B, Mistral-7B-Instruct-v0.2 and gpt-3.5-turbo-0125333https://openai.com/api/ for Llama3, Mistral and GPT-3.5 respectively.
3.1.4 Example Selection
@@
3.1.5 Prompts
We compare ERR with Vanilla across different datasets and LLMs. Given that these tasks usually involve reasoning, on which chain-of-thought (CoT) is commonly utilized Wei et al. (2022), we experiment both without CoT ("w/o CoT", which are identical to the original templates) and with CoT ("w/ CoT"). This allows us to examine both the orthogonality and compatibility with CoT of ERR , as well as assess its performance across various models and tasks. Specifically, we evaluate Vanilla (w/o CoT), Vanilla (w/ CoT), ERR (w/o CoT), and ERR (w/ CoT). Due to space constraints, examples of prompt templates discussed in this section are provided in Appendix D.
Figure 7: Results of the four types of tasks on GPT-3.5.
@@ 3.2 Results of Small-scale Models
Results across all nine datasets and three small-scale models (Alpaca, Llama3 and Mistral) are illustrated in Figure 6.
The results demonstrate that ERR (w/ CoT), achieved by integrating CoT with our proposed prompt framework, either significantly outperforms or matches Vanilla (w/ CoT) in 25 out of 27 experiments (covering 9 datasets and 3 models). The exceptions are Alpaca on the Sports dataset and Mistral on the AQuA dataset, where ERR (w/ CoT) shows somewhat lower performance compared to Vanilla (w/ CoT). When CoT is not incorporated, ERR generally performs much better than or on par with Vanilla, except for the Sports dataset with Llama3, where ERR performs a little poorer.
Surprisingly, ERR (w/o CoT) sometimes even surpasses Vanilla (w/ CoT), suggesting that the ERR framework alone can offer more improvements than CoT. This highlights the value of ERR and reaffirms that the format plays a crucial role in enhancing LLMs’ problem-solving capabilities. In terms of models, the performance of ERR on Alpaca is far less impressive than on Llama3 and Mistral, which may be because Alpaca has strong instruction-following capabilities and is more robust to different prompts.
@@
We also experiment with Llama2 Touvron et al. (2023b) and the results are in Appendix E. The overall trend is consistent with Llama3.
@@
As shown in Figure 7, ERR performs similarly to Vanilla across every dataset using GPT-3.5. Although the ERR format does not bring significant improvement to these tasks with GPT-3.5 and Alpaca (as shown in Figure 6), the fact remains that the incorrect or meaningless example-level descriptions caused by random nouns do not have much negative impact on GPT-3.5, a sufficiently powerful model, or Alpaca, which has strong instruction-following capabilities. In some cases, it even slightly improves performance (e.g., ERR (w/ CoT) outperforms Vanilla (w/ CoT) on AQuA and Known Unknowns). In other words, LLMs might not care what users actually say to describe the provided examples while they are more sensitive to the format of prompts, which is in line with our findings obtained from MT.
@@ 3.4 Discussion
Based on the experiments conducted on both small-scale and large-scale models, we can conclude that ERR is a simple yet practical and universal prompt framework. It can enhance problem-solving capabilities in small models and be applied to large models without the risk of performance degradation due to the meaningless noise within it. In other words, there might be less need to meticulously select examples or design detailed descriptions. Instead, you can uniformly and efficiently apply ERR to various tasks with different models.
@@ 4 Related Work
Rubin et al. Rubin et al. (2022) suggest that LLMs’ ICL performance strongly depends on the selection of in-context examples. In consequence, many works have been trying to explore ways of selecting better in-context examples in recent years. Li et al. Li et al. (2023) train a unified in-context example retriever across a wide range of tasks. Ye et al. Ye et al. (2023) select examples based on both relevance and diversity, with the help of determinantal point processes. Agrawal et al. Agrawal et al. (2023) ensure n-gram coverage to select better examples for MT. Kumar et al. Kumar et al. (2023) train an in-context example scorer for MT based on several features. Tang et al. Tang et al. (2024) propose to combine both word-level and syntax-level coverage when selecting examples for MT.
@@
With the popularity of ICL, there have been numerous studies on analyzing the mechanism of ICL. One stream of these studies focuses on explaining the essence of ICL, relating ICL to gradient descent Von Oswald et al. (2023), implicit Bayesian inference Xie et al. (2022), induction heads completing token sequences based on similar context Olsson et al. (2022), generation maintaining coherency Sia and Duh (2023), creation of task vectors based on in-context examples Hendel et al. (2023), etc. The other stream focuses on the role of in-context examples, especially labels of these examples. Min et al. Min et al. (2022) find that ground truth labels are not necessary and LLMs perform fairly well even with random labels. Wang et al. Wang et al. (2023) find that label words play the role of anchors that aggregating information of the whole examples and serve as a reference for LLMs’ final predictions. Wei et al. Wei et al. (2023) find that larger language models can override semantic priors and learn from in-context examples with flipped labels or semantically-unrelated labels.
@@ 5 Conclusion
In this work, we analyze the effect of descriptive instructions in prompts during ICL and propose an Ensemble prompt framework describing the properties of in-context examples selected by different methods. Experimental results on MT indicate that while LLMs are sensitive to prompt formats, they might not care the actual meaning of the descriptions and the framework improves LLMs’ performance even with meaningless descriptions compared with the conventional prompt. We further apply the Ensemble framework to four other NLP tasks and find that it achieves promising results, especially on small-scale models. These results suggest that rather than working hard on well-designed descriptions, making use of a proper prompt format would be more effective and efficient.
Limitations
First, since there are so many open-source LLMs in the world nowadays, it is impossible to experiment with all existing models and thus our work only employ several commonly-used LLMs. Second, since we do not have access to the pre-training or post-training process of LLMs (either open-source or close-source), our analysis of the mechanism of ICL could be somewhat superficial. The behavior of LLMs can be highly subject to their training data, which we have no access to. Lastly, although we reveal that ERR is a superior prompt format for several models, it could still be a local optimum and how to effectively search for a best prompt format for different models and tasks is still under-explored, which we leave for future work.
Appendix A Prompt Templates of MT Experiments
Figure 8: Template and example of Vanilla.
Figure 9: Template and example of Ensemble (Word + Syntax).
Figure 10: Template and example of Ensemble (Syntax + Word).
Figure 11: Template and example of Different Ensemble (Word + Syntax).
Figure 12: Template and example of Ensemble (Word + Semantics).
Figure 13: Template and example of Ensemble (Random + Random).
Figure 14: Template and example of Single (Random).
Figure 15: Template and example of Single (Example).
Figure 16: Template and example of Vanilla (Translate).
A.1 Main Experiments
Templates and prompt examples of Vanilla, Ensemble (Word + Syntax), Ensemble (Syntax + Word), Different Ensemble (Word + Syntax), Ensemble (Word + Semantics) and Ensemble (Random + Random) are shown in Figure 8-13.
A.2 Ablations
Templates and prompt examples of Ensemble (Random + Random), Single (Random), Single (Example), Vanilla (Translate) are shown in Figure 13-16.
Appendix B Attention Weights on Alpaca
Figure 17: Attention weights (×1e-4) on Alpaca of all 32 layers with BM25 + Polynomial examples. EWS and ERR denotes Ensemble (Word + Syntax) and Ensemble (Random + Random) respectively.
Figure 17 presents the attention weights on Alpaca. For example-to-noun attention weights, ERR is close to EWS. For target-to-noun attention weights, EWS is higher in shallow layers but falls behind ERR in deeper layers, especially in the last layer. This demonstrates that Alpaca might pay more attention to the meaningful words ("word" and "syntax") when understanding the context in shallow layers but gradually forgets them when it comes to generation in the deeper layers. In short, EWS performs no higher than ERR in most cases.
Appendix C Dataset Details for Reasoning Tasks
We list the details of splitting training set (example database) and test set for our conducted reasoning tasks, covering four types and nine datasets. We set random seed for all possible shuffling and sampling operations to $42$. Note that we experiment with 4-shot for all datasets.
C.1 Datasets Fetched from Exclusive Source
  • CSQA Talmor et al. (2019): https://www.tau-nlp.org/commonsenseqa. We follow the official split and select the training set as our example database and the dev set as our test set. Because the training set itself is randomly divided from the whole dataset, we directly select examples from it in order.
  • GSM8K Cobbe et al. (2021): https://github.com/openai/grade-school-math. We select the test.jsonl as our test set and the train.jsonl as our example database and randomly sample four examples from it.
  • AQuA Ling et al. (2017): https://github.com/google-deepmind/AQuA. We select the test.json as our test set. Since the original training set is relatively large, for simplicity, we directly copy the four examples listed in the supplementary materials of Wei et al. (2022) and we ensure that these four examples do not appear in the test set.
C.2 Datasets Fetched from The Big-bench
For StrategyQA Geva et al. (2021), Date, Sports, Logical Fallacy, Three Objects, and Known Unknowns, we fetched them from the Big-bench Srivastava et al. (2023). Each of them has a task.json. We randomly shuffle the task.json and split it to a training set and a test set. Then we select examples from the training set in order.
Specifically, the principle for splitting the training and test sets is as follows: If the total number of samples exceeds 1,012 a lot, we retain 1,012 samples as the test set and use the remainder as the training set. Otherwise, we select four examples for the training set and use the rest for testing. For the Sports and Logical Fallacy datasets, which have only two possible answers (similar to binary classification), we first separate the positive and negative examples, shuffle them individually, and then construct the test set and training set. The test set is composed of an equal number of positive and negative examples, with the remaining samples used as the training set.
Appendix D Prompts for Reasoning Tasks Used in this Work
Figure 18: Prompt for CSQA.
Figure 19: Prompt for StrategyQA.
Figure 20: Prompt for Date.
Figure 21: Prompt for Sports.
Figure 22: Prompt for Logical Fallacy.
Figure 23: Prompt for Three Objects.
Figure 24: Prompt for Known Unknowns.
Figure 25: Prompt for GSM8K. For this dataset, we let LLMs first generate reasoning and then answer under the "w/ CoT" setting.
Figure 26: Prompt for AQuA.
Figure 18-26 show the examples of ERR (w/ CoT) prompt for respective datasets. Some tasks contain Answer Choices. In order to save space, the blank lines between the options are replaced with spaces in those figures. Each Figure has grey text for reasoning, cyan text for the parts of ERR that are unique to Vanilla, and italic words in the cyan text representing random nouns. Therefore, deleting the grey text gives ERR (w/o CoT), keeping the grey text but deleting the cyan text gives Vanilla (w/ CoT), and deleting both the cyan and grey text gives Vanilla (w/o CoT). [^4] When changing ”w/ CoT” to ”w/o CoT”, you may also need to replace ”So the answer is …” with ”The answer is …” for syntactical reasons. The reasoning is generated by ChatGPT  [^5] https://chatgpt.com/. Note the ChatGPT is not the same as GPT-3.5 we used for experiments.
Appendix E Results of Llama2
Figure 27: Results of Llama2 on the nine datasets.
Results of Llama2-7B-chat-hf Touvron et al. (2023b) on the nine datasets are presented in Figure 27. While ERR outperforms Vanilla with Llama2 across most datasets, its performance on Logical Fallacy and Sports is notably poor. Llama2 almost always responds with confused emojis for Logical Fallacy questions and outputs questions like "plausible or implausible?" for Sports, leading to predominantly incorrect answers. Further investigation into these issues is left for future work.
Appendix F Computational Details
F.1 Hardware
Inference of LLMs runs on an NVIDIA A40 GPU (with memory of 48 GB). Other experiments run on Intel® Xeon® Gold 6348 CPU (with memory of 256 GB).
F.2 Software
Our OS: Ubuntu 20.04.6 LTS. Our code: Python only. Libraries and packages are specified in the source code.
Appendix G Licenses
Artifact License
XGLM MIT
Alpaca Apache-2.0
Llama Llama Community License Agreement
Mistral Apache-2.0
COMET Apache-2.0
FLORES-101 CC-BY-SA-4.0
Europarl Unknown
ParaCrawl CC0
CSQA CC-BY-SA-4.0
StrategyQA MIT
BIG-bench Apache-2.0
GSM8K MIT
AQuA Apache-2.0
Table 4: Licenses of scientific artifacts we use.
The licenses of the scientific artifacts we use are shown in Table 4.